Aaleighacruz1:

Pyramid A is a square pyramid with a base side length of 14 inches and a height of 8 inches. Pyramid B has a volume of 3,136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Provide an explanation and proof for your answer to receive full credit. (10 points)

surjithayer:

find the volume of Pyramid A. volume of pyramid=base area * height \[area ~of ~square =side ^2\] then divide the volume of Pyramid B by volume of Pyramid A

Aaleighacruz1:

224

surjithayer:

can you solve it?

surjithayer:

how you found it? can you show me?

Aaleighacruz1:

I divided the volume of Pyramid B by volume of Pyramid A

surjithayer:

base area=14^2=196 sq. in volume of Pyramid A=196*8= 1568 ~cubic inches now can you divide/

Aaleighacruz1:

divide what?

surjithayer:

divide the volume of Pyramid B with volume of Pyramid A

Aaleighacruz1:

8?

surjithayer:

3136 /1568=2

surjithayer:

8 is not correct.

Aaleighacruz1:

so 2?

surjithayer:

yes

Aaleighacruz1:

so 2 times bigger?

surjithayer:

correct.

Aaleighacruz1:

thank you!

surjithayer:

yw